I need to solve the equation:
5-x 13
___=__
17 34

Answer by stanbon(75887)

solve the equation:
5-x 13
___=__
17 34
==============
Multiply both sides by 34 to get:
2(5-x)=13
10-2x=13
2x=-3
x=-3/2
=======
cheers,
Stan H.